OpenStudy (amilapsn):

when does the minute hand and hour hand come one on one after 12 o' clock?

Parth (parthkohli):

It's around 1:05. You can use math to get a more precise answer.

OpenStudy (amilapsn):

thnk you

Parth (parthkohli):

Let's say it's 1 hr and \(x\) min. What does it mean when two hands coincide? How do we put it mathematically? It means that we can make an equation. What do we equate? The angles. We equate the angles. At 1 hour and \(x\) minutes, the angle created by the hour-hand wrt 12 is:\[30^{\circ}\left( 1+\dfrac{x}{60}\right)\]and that created by the minute hand wrt 12 is:\[360^{\circ }\cdot\dfrac{x}{60}\]Equate those two and you get\[x = \frac{60}{11} = 5.4545\cdots\]

Parth (parthkohli):

Therefore, the time is 1 hour and 5.4545... minutes. Or around 1:05:27.

ganeshie8 (ganeshie8):

is it possible to model the situation wid vectors

OpenStudy (dan815):

ye

OpenStudy (dan815):

we know both angular velocity so

OpenStudy (dan815):

u can write position vectors for both and equate them together

Parth (parthkohli):

oh wow

ganeshie8 (ganeshie8):

t is in hours : minute hand : <cos(2pi*60*t), sin(2pi*60*t)> hour hand : <cos(2pi*1*5m), sin(2pi*1*t)> how to solve

OpenStudy (dan815):

i think its better to write it out as arc length wrt to time

OpenStudy (dan815):

and equate the 2 sinusoids and u shud get 12 solutions which is easy to see from the clocks and the periods that are forming

Parth (parthkohli):

\[\rm \omega_{second}= 2\pi / T = 2\pi ~rad / 60 sec = \pi/30 ~rad ~sec^{-1}\]\[\rm \omega_{hour} = 2\pi / T = 2\pi ~rad / (60\times 60)sec = \pi /1800 ~ rad~sec^{-1}\]We equate the positions now? (The position of second hand can be a multiple of that of the hour hand.)

OpenStudy (dan815):